OpenStudy (anonymous):

he difference between Sheila and Robbie's ages is 12. The sum of their ages is 42. If Sheila is the younger of the two, how old is Robbie?

OpenStudy (anonymous):

the

OpenStudy (anonymous):

Total: 42 Difference: 12 42/2 = 21 21 - 12 = 9 That means Sheila is 9, since she is the youngest of the two. Robbie is 21.

OpenStudy (anonymous):

there is no 21 theres only 15,27.29,32

OpenStudy (anonymous):

Lol oh! Well, what do you think the answer is?

OpenStudy (anonymous):

a-b=12 a+b=42 add 2a=54 a=27 so b=15

OpenStudy (anonymous):

15 rounds to 21